The Matsui Lubricants plant uses the FIFO method to account for its work-in-process inventories. The accounting records show the following information for a particular day:

    

  Beginning WIP inventory
     Direct materials $ 1,926
     Conversion costs 514
  Current period costs
     Direct materials 22,490
     Conversion costs 13,414

  

Quantity information is obtained from the manufacturing records and includes the following:

  

  Beginning inventory 610 units (70% complete as to materials,
50% complete as to conversion)
  Current period units started 4,100 units
  Ending inventory 1,400 units (40% complete as to materials,
10% complete as to conversion)

  

Required:

Compute the cost per equivalent unit for direct materials and for conversion costs using the FIFO method. (Round your answers to 2 decimal places.)
Answer #1

Equivalent unit :

Whole unit % EUP-material % EUP-conversion
Beginning work in process 610 30% 183 50% 305
Unit started and completed 2700 100% 2700 100% 2700
Ending work in process 1400 40% 560 10% 140
Equivalent unit of production 4710 3443 3145

Cost per equivalent unit of material = 22490/3443 = $6.53

Cost per equivalent unit of conversion = 13414/3145 = $4.27